OpenStudy (anonymous):

If f(x) = −x − 1 and g(x) = 4x − 1, find f[g(−2)] I think I know the answer, but i'm not positive!

OpenStudy (anonymous):

@Luis_Rivera @RadEn @ganeshie8 can you please help me?

OpenStudy (raden):

g(x) = 4x − 1 g(-2) = 4(-2) - 1 = ...

OpenStudy (anonymous):

-9?

OpenStudy (raden):

yup. correct now f[g(−2)] = f(-9) subitute x =-9 into f(x) = -x-1 so, f(-9) = -(-9) - 1 = ...

OpenStudy (anonymous):

so the answer is -10?

OpenStudy (raden):

-(-9) - 1 = 9 - 1 = 8

OpenStudy (anonymous):

that was my original answer :) thank you!

OpenStudy (raden):

welcome